OpenStudy (cookiemonster18):

Solve the Equation: 2x - 6 = 2(x - 8) + 10 A. x=2 B. x=8 C. Infinitely many solutions D. No solution

OpenStudy (calculusxy):

2x - 6 = 2(x - 8) + 10 2x - 6 = 2x - 16 + 10 2x - 6 = 2x - 6 these two equal to the same thing, so it's infinite number of solutions

OpenStudy (anonymous):

2(8)-6 16-6=10 2(8-8)+10 2(0)+10 0+10=10

OpenStudy (cookiemonster18):

is it C.

OpenStudy (calculusxy):

it basically means that no matter what number you put into that equation, it will always be equal to each other

OpenStudy (cookiemonster18):

oh okay @calculusxy but I was thinking it was D because I had tiger algebra solve it and it said 0=0 but now I think it's C. am I correct,that it's C?

OpenStudy (calculusxy):

i think it's C.